Suppose that 8 years ago, you purchased 284 shares of stock in a corporation. Between then and now, the stock had a 2:1 split and a 4:1 split. Today, each share sells for $11. If selling all of your shares today would give your investment an annual rate of return of 6%, what was the price per share when you made the purchase 8 years ago? Round your answer to the nearest dollar.
Answer #1
Orignial share purchased 284
Number of share after 2:1 split =284*2 568
Number of share after 4:1 split =568*4 2272
Number of share outstanding 2272
Value of share = 2272*11 24992
Annual return over 8 year period 6%
Intial cost = 24992/(1.06)^8 15680.29
Orignial share = 284
Price per share = 15680.29/284             55
Price per share 8 year ago = $55